how far dose a car go a minute if it going 20 kilometers an hour

Answer by Maths68(1474)

1 hour............20km
60 minutes........20km
60/60 minutes.....20/60 km
1 minute.........1/3 km

Or
Distance = ?
Time = 1 minute
Speed = 20km/h
Distance = Speed * time
Distance = 20/hour * time
Distance = 20/60 minutes * 1 minute
Distance = 1/3 * 1
Distance = 1/3 km